# TimeBridge: Non-Stationarity Matters for Long-term Time Series Forecasting

### v2

@@ 5 Ablation Studies
Case Integrated Attention Cointegrated Attention Weather Solar Electricity Traffic
@@
Case Integrated Attention Cointegrated Attention Weather Solar Electricity Traffic
@@
Case Integrated Attention Cointegrated Attention Weather Solar Electricity Traffic
@@
6 Non-stationarity and Dependency Modeling Analysis
Intra-variate Modeling. As shown in Fig. 4a, when non-stationarity is retained, the attention map in the temporal dimension diverges, with the model focusing on multiple patches across a broader time span. However, after removing non-stationarity, the attention map becomes more concentrated on adjacent time steps, aligning with the causal nature of time series, where closer time steps are usually more correlated. Non-stationarity may cause the model to mistake distant similarities for causality. By eliminating non-stationarity, the model better captures short-term variations and local dependencies, enhancing its robustness and interpretability in handling complex time series data.
Inter-variate Modeling. Fig. 4b shows that removing non-stationarity narrows the model’s attention to a few inter-variate dependencies, while retaining non-stationarity enables the capture of more diverse and richer relationships. Non-stationary sequences help the model identify cointegration, revealing hidden equilibrium mechanisms in multivariate time series. Preserving non-stationarity enhances the model’s ability to express complex inter-variate dependencies. Additionally, Fig. 4c shows the impact of different patch downsampling rates on performance. For datasets with more channels and stronger cointegration (e.g., Solar and Traffic), increasing downsampled patches initially improves predictions by preserving long-term features. However, too much downsampling adds computational cost and negatively affects smaller-channel datasets (e.g., Weather), so we carefully balanced downsampling rates based on dataset characteristics, as detailed in Tab. 9.
Real Case of Weather Forecast. Given the strong interrelationships between weather variables, we analyze temperature $T$ and dew point temperature $T_{\text{dew}}$ from the Weather dataset. Dew point measures atmospheric moisture and is typically closely linked to temperature. Without external influences, such as water vapor or heat sources, the difference between temperature and dew point is minimal, showing long-term cointegration. However, temperature tends to exhibit more short-term fluctuations due to external factors like sunlight and weather systems. As shown in Fig. 5, the results demonstrate that spurious regressions can only be avoided by eliminating non-stationarity during short-term modeling, while preserving it during long-term dependency modeling to capture the underlying cointegration between variables.
@@ Appendix C Datasets
We conduct extensive experiments on eight widely-used time series datasets for long-term forecasting. Additionally, we use the PeMS datasets for short-term forecasting and the CSI 500 and S&P 500 indices for financial forecasting. We report the statistics in Tab. 8. Detailed descriptions of these datasets are as follows:
@@
    Solar-Energy dataset (Lai et al., 2018) contains solar power production data recorded every 10 minutes throughout 2006 from 137 photovoltaic (PV) plants in Alabama.
@@ Appendix D Implementation Details
All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on two NVIDIA RTX 3090 24GB GPUs. We use the Adam optimizer (Kingma, 2014). The batch size is set to 16 for the Electricity and Traffic datasets, and 32 for all other datasets. All models are trained for 10 epochs. Tab. 9 provides detailed hyperparameter settings for each dataset. For the four ETT datasets, the relatively small number of channels results in less pronounced long-term cointegration relationships, as evidenced by the low EG test results in Tab. 8. Therefore, we focus on modeling short-term intra-variate variations only.
Num. of Integrated Num. of Cointegrated $N$ $M$ lr d_model d_ff
ETTh1 $2$ 00 $30$- $1e$-$4$ $128$ $256$
ETTh2 $2$ 00 $30$- $1e$-$4$ $64$ $128$
ETTm1 $2$ 00 $30$- $1e$-$4$ $32$ $128$
ETTm2 $2$ 00 $30$- $1e$-$4$ $32$ $64$
Weather $1$ $1$ $30$ $12$ $1e$-$4$ $128$ $128$
Solar $1$ $1$ $30$ $12$ $5e$-$4$ $128$ 128
Electricity $2$ $1$ $30$ $4$ $5e$-$4$ $512$ 512
Traffic $2$ $2$ $30$ $4$ $1e$-$3$ $512$ 512
@@ E.1 Main Experiments
Tab. 10 and Tab. 11 present the full results for long-term forecasting, including both the original results from their respective papers and those obtained through hyperparameter search. The hyperparameter search process involved exploring input lengths $I\in\{96,192,336,512,720\}$, learning rates from $10^{-5}$ to $0.05$, encoder layers from 1 to 5, $d_{model}$ values from 16 to 512, and training epochs from 10 to 100. In both settings, TimeBridge consistently achieved the best performance, demonstrating its effectiveness and robustness. Additionally, for financial forecasting, we included three additional strong baselines: ALSTM (Qin et al., 2017), GRU (Chung et al., 2014), and TRA (Lin et al., 2021). The results in Tab. 12 show that TimeBridge continues to outperform these methods, further validating its superiority.
Models TimeBridge iTransformer PDF TimeMixer PatchTST Crossformer FEDformer ModernTCN MICN TimesNet DLinear

### v3

@@ 5 Ablation Studies
Case Integrated Cointegrated Weather Solar Electricity Traffic
Attention Attention
@@
Case Integrated Cointegrated Weather Solar Electricity Traffic
Attention Attention
@@
Case Integrated Cointegrated Weather Solar Electricity Traffic
Attention Attention
@@
6 Non-stationarity and Dependency Modeling Analysis
Intra-variate Modeling. As shown in Figure 4a, when non-stationarity is retained, the attention map in the temporal dimension diverges, with the model focusing on multiple patches across a broader time span. However, after removing non-stationarity, the attention map becomes more concentrated on adjacent time steps, aligning with the causal nature of time series, where closer time steps are usually more correlated. Non-stationarity may cause the model to mistake distant similarities for causality. By eliminating non-stationarity, the model better captures short-term variations and local dependencies, enhancing its robustness and interpretability in handling complex time series data.
Inter-variate Modeling. Figure 4b shows that removing non-stationarity narrows the model’s attention to a few inter-variate dependencies, while retaining non-stationarity enables the capture of more diverse and richer relationships. Non-stationary sequences help the model identify cointegration, revealing hidden equilibrium mechanisms in multivariate time series. Preserving non-stationarity enhances the model’s ability to express complex inter-variate dependencies. Additionally, Figure 4c shows the impact of different patch downsampling rates on performance. For datasets with more channels and stronger cointegration (e.g., Solar and Traffic), increasing downsampled patches initially improves predictions by preserving long-term features. However, too much downsampling adds computational cost and negatively affects smaller-channel datasets (e.g., Weather), so we carefully balanced downsampling rates based on dataset characteristics, as detailed in Tab. 8.
Real Case of Weather Forecast. Given the strong interrelationships between weather variables, we analyze temperature $T$ and dew point temperature $T_{\text{dew}}$ from the Weather dataset. Dew point measures atmospheric moisture and is typically closely linked to temperature. Without external influences, such as water vapor or heat sources, the difference between temperature and dew point is minimal, showing long-term cointegration. However, temperature tends to exhibit more short-term fluctuations due to external factors like sunlight and weather systems. As shown in Figure 5, the results demonstrate that spurious regressions can only be avoided by eliminating non-stationarity during short-term modeling, while preserving it during long-term dependency modeling to capture the underlying cointegration between variables.
@@ 7 Conclusion
Impact Statement
Time series forecasting is fundamental in various real-world applications, including finance, weather prediction, and traffic management. Our work introduces an innovative approach to handling non-stationarity, improving both short-term and long-term dependency modeling. The datasets used in this study are publicly available, ensuring reproducibility and transparency. We do not foresee any ethical concerns associated with this research. Our contributions primarily advance the field of time series forecasting, with potential benefits across multiple domains.
@@ Appendix C Datasets
We conduct extensive experiments on several widely-used time series datasets for long-term forecasting. Additionally, we use the PeMS datasets for short-term forecasting and the CSI 500 and S&P 500 indices for financial forecasting. We report the statistics in Tab. 7. Detailed descriptions of these datasets are as follows:
@@
    Solar-Energy dataset (Liu et al., 2024a) contains solar power production data recorded every 10 minutes throughout 2006 from 137 photovoltaic (PV) plants in Alabama.
  6. (6)
    Climate dataset (Godahewa et al., 2021) records regional precipitation data across 1,763 stations in Australia.
@@ Appendix D Implementation Details
All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on two NVIDIA RTX 3090 24GB GPUs. We use the Adam optimizer (Kingma, 2014). The batch size is set to 16 for the Electricity and Traffic datasets, and 32 for all other datasets. All models are trained for 10 epochs. Tab. 8 provides detailed hyperparameter settings for each dataset. For the four ETT datasets, the relatively small number of channels results in less pronounced long-term cointegration relationships, as evidenced by the low EG test results in Tab. 7. Therefore, we focus on modeling short-term intra-variate variations only. To improve training stability, we adopt a hybrid MAE loss that operates in both the time and frequency domains. Compared to MSE, which amplifies differences, MAE provides a more stable optimization process. Additionally, the frequency-domain loss mitigates label autocorrelation, making training more effective. The loss function is defined as:
$L_{t}=\frac{1}{N}\sum_{i=1}^{N}|\mathbf{X}_{i}-\hat{\mathbf{X}}_{i}|$ $,\quad L_{f}=\frac{1}{N}\sum_{i=1}^{N}|\text{FFT}(\mathbf{X}_{i})-\text{FFT}(\hat{\mathbf{X}}_{i})|,$
$L$ $=(1-\alpha)\times L_{t}+\alpha\times L_{f},$
where FFT denotes the Fast Fourier Transform. $\alpha$ is the hyperparameter.
Num. of Integrated Num. of Cointegrated $N$ $M$ lr d_model d_ff $\alpha$
ETTh1 $2$ 00 $30$ $16$ $1e$-$4$ $128$ $128$ 0.35
ETTh2 $2$ 00 $12$ $12$ $1e$-$4$ $128$ $128$ 0.8
ETTm1 $2$ 00 $15$ $6$ $2e$-$4$ $64$ $128$ 0.8
ETTm2 $2$ 00 $30$ $12$ $2e$-$4$ $32$ $64$ 0.35
Weather $1$ $1$ $30$ $12$ $1e$-$4$ $128$ $128$ 0.1
Solar $1$ $1$ $30$ $12$ $5e$-$4$ $128$ $128$ 0.05
Electricity $1$ $2$ $30$ $4$ $5e$-$4$ $512$ $512$ 0.2
Traffic $2$ $2$ $30$ $4$ $1e$-$3$ $512$ $512$ 0.8
Climate $2$ $2$ $30$ $4$ $1e$-$3$ $768$ $768$ 0.2
@@ E.1 Main Experiments
Tab. 9 present the full results for long-term forecasting through hyperparameter search. The hyperparameter search process involved exploring input lengths $I\in\{96,192,336,512,720\}$, learning rates from $10^{-5}$ to $0.05$, encoder layers from 1 to 5, $d_{model}$ values from 16 to 512, and training epochs from 10 to 100. In both settings, TimeBridge consistently achieved the best performance, demonstrating its effectiveness and robustness.
Furthermore, on the large Climate dataset (Godahewa et al., 2021), which contains 1,763 channels, the prediction task becomes particularly challenging due to the more complex inter-channel dependencies and higher levels of noise in the data. Despite these challenges, TimeBridge maintains its superior performance over competing models.
Additionally, for financial forecasting, we included three additional strong baselines: ALSTM (Qin et al., 2017), GRU (Chung et al., 2014), and TRA (Lin et al., 2021). The results in Tab. 10 show that TimeBridge continues to outperform these methods, further validating its superiority.
Models TimeBridge iTransformer
DeformableTST
TimeMixer PatchTST Crossformer FEDformer ModernTCN MICN TimesNet DLinear